# ChatENV: An Interactive Vision-Language Model for Sensor-Guided Environmental Monitoring and Scenario Simulation

### v1

@@ Abstract
Understanding environmental changes from aerial imagery is vital for climate resilience, urban planning, and ecosystem monitoring. Yet, current vision language models (VLMs) overlook causal signals from environmental sensors, rely on single-source captions prone to stylistic bias, and lack interactive scenario-based reasoning. We present ChatENV, the first interactive VLM that jointly reasons over satellite image pairs and real-world sensor data. Our framework: (i) creates a 177k-image dataset forming 152k temporal pairs across 62 land-use classes in 197 countries with rich sensor metadata (e.g., temperature, PM10, CO); (ii) annotates data using GPT-4o and Gemini 2.0 for stylistic and semantic diversity; and (iii) fine-tunes Qwen-2.5-VL using efficient Low-Rank Adaptation (LoRA) adapters for chat purposes. ChatENV achieves strong performance in temporal and “what-if” reasoning (e.g., BERT-F1 0.902) and rivals or outperforms state-of-the-art temporal models, while supporting interactive scenario-based analysis. This positions ChatENV as a powerful tool for grounded, sensor-aware environmental monitoring. [^1] The annotated data and code are available on github through: https://github.com/HosamGen/ChatENV/
Index Terms:
Environmental Monitoring, Vision-Language Models, Scenario Prediction, Remote Sensing.
@@ I Introduction
Monitoring and understanding environmental dynamics are top priorities for addressing pressing issues like climate change, urban growth, and habitat degradation. Remote sensing, particularly through the use of aerial and satellite imaging, offers extensible monitoring of Earth’s surface, enabling real-time detection of changes in land cover, vegetation, urban areas, and other environmental indicators [1, 2]. Recently, Vision-Language Models (VLMs) have come into the spotlight for fusing visual and text data, enabling enhanced automation of complex tasks in remote sensing, ranging from change detection, captioning, and multimodal reasoning [3, 4]. In spite of this, several key challenges remain:
(1) Lack of environmental context, as most VLMs rely solely on images and ignore essential sensor data crucial for understanding and observing geographical changes [3]; (2) Single-source annotation bias, where using a single language model for data annotation limits linguistic diversity and introduces bias [5]; (3) Limited spatial and semantic diversity, since widely used datasets like LEVIR-CD [6] and RSICap [7] lack the geographic scope and object-class variety necessary for robust modeling [8]; and (4) Lack of interactive reasoning, with current VLMs offering no "what-if" interfaces to support planning and hypothetical decision-making [3].
@@
  • SOTA Multimodal Dataset. We release the largest dataset of its kind, containing over 177k satellite images, divided into 152k time-separated image pairs, across 62 object classes and 197 countries. Each pair is accompanied by rich change captions generated by a dual-model pipeline (GPT-4o and Gemini 2.0), adding linguistic variety, context richness, and improved generalization. All test captions are manually reviewed to ensure accuracy, providing a strong benchmark for environmental change detection tasks.
@@ II Related Work
Benchmark Datasets: Several datasets support geospatial vision–language tasks. For captioning, RSICD [8], NWPU-Captions [9], and RSICap [7] offer image-text pairs at varying scales. RSVQA [10] and BigEarthNet-based extensions [11] focus on spatial Visual Question Answering (VQA), while LEVIR-CD [12] and LEVIR-CC [13] provide Change Detection (CD) labels and natural language descriptions. ChangeCLIP [14] benchmarks semantic CD, and SpaceNet [15] provides long-term urban imagery but lacks text annotations. However, most existing datasets lack temporal variety, environmental sensor data, or support for open-ended user queries. There remains a need for benchmarks that support contextual understanding, multimodal learning, and interactive reasoning.
Vision-Language Models: Recent VLMs such as Pixtral [16], DeepSeek-VL [17], Janus Pro [18], Phi-4-Multimodal [19], and LLaMA-3-Vision [20] demonstrate strong performance on tasks like captioning, visual reasoning, and instruction following. These models integrate vision encoders directly into Large Language Models (LLMs) and are trained on large-scale image–text datasets. While not designed for remote sensing, their scalability and reasoning capabilities make them promising candidates for domain adaptation.
VLMs for Remote Sensing: Domain-specific adaptations of VLMs have emerged to tackle geospatial tasks. For example, RSGPT [7] fine-tunes Vicuna on RSICap for aerial captions. RS-LLaVA [21] and EarthGPT [22, 23] use CLIP/ViT backbones with LLaMA-based LLMs for captioning and VQA, incorporating multi-sensor data. RemoteCLIP and GeoChat [24, 25] build on large image–text corpora but lack temporal modeling. GeoLLaVA [26] uses video-based fine-tuning for change detection, though it still lacks annotation diversity and interactivity.
Captioning and VQA in Aerial Imagery: Initial work on remote sensing VQA began with RSVQA [10], which used OpenStreetMap for spatial questions, and was extended by [11] using BigEarthNet. These efforts rely on synthetic prompts and lack environmental data. In captioning, RSICD introduced a CNN–RNN pipeline, with NWPU-Captions adding attention mechanisms, and more recent work [27] exploring metadata-aware pretraining but is still limited to static imagery.
@@
RSGPT [7]
RS-LLaVA [22]
EarthGPT [23]
RemoteCLIP [24]
GeoChat [25]
GeoLLaVA [26]
Tree-GPT [28]
ChangeCLIP [14]
CD-VQA [11]
TEOChat [29]
@@
Interactive Reasoning for Environmental Monitoring: Conversational VLMs like, RS-LLaVA, and EarthGPT enable multi-turn Q&A but lack predictive reasoning. Tree-GPT [28] adds geospatial tools for command-based interaction but cannot simulate hypothetical, counterfactual scenarios. Finally, TEOChat [29] supports scene monitoring by asking questions about objects that change in an image or disasters affecting a location, but it does not incorporate environmental reasoning based on temperature or other meteorological metrics.
@@ III Data Collection and Pre-processing
Construct Temporal Pairs from fMoW Dataset. fMoW RGB dataset [30] consists of 363,571363{,}571 train images and 53,04153{,}041 validation images covering years $2002$ to $2017$ globally. With spatial resolution around $0.3$ mm, it provides increased accuracy for fine-grained change detection between $62$ classes [31]. We utilized the dataset’s metadata to pair images based on location while maintaining a minimum temporal gap of $12$ months, as preliminary inspection of the dataset indicated that visible and detectable changes in the scenes typically emerge after approximately one year. Upon filtering and removing duplicates, the final training set has $177k$ images, corresponding to $152k$ pairs, and the test set consists of $14k$ images, which correspond to $10k$ pairs. Table II and Figure 4 provide an overview and a visual distribution of the 62 classes over 197 countries, respectively.
@@
We used fMoW Dataset metadata, which contains geospatial locations, country codes, time zones, and timestamps, to fetch weather and atmospheric emissions data through third-party APIs. For weather features, the VisualCrossing API [32] fetched a large set of fields, like temperature, dew point, humidity, and more, by using the capture location and timestamp of each image. Likewise, emissions data were collected using Open-Meteo [33] for European samples and OpenAQ [34] for non-European samples, yielding parameters like $pm2.5$, COCO, $NO_{2}$, and ozone. We concentrated on features that appeared most frequently and for which our analyses would be most relevant, and cross-referenced country codes from the fMoW dataset against the country name returned by the APIs to ensure temporal alignment and season-specific context. Features with limited relevance or significant missing data were discarded, while those lacking emissions data were retained to preserve category and geographic diversity. This fusion-mapping of environmental metadata and the dataset’s global, multi-year coverage allowed us to capture temporal, climatic, and atmospheric changes with minimal missing information, supporting downstream modeling work.
@@
We produced descriptive annotations to capture the most important features of each image and summarize temporal variation between matched samples. Two SOTA LLMs, GPT-4o and Gemini 2.0 Flash Thinking, were used to extend generalization, mitigate stylistic bias, and increase robustness. All prompts included season, semantic category, and environmental sensor data, following a consistent prompting strategy for each model. The dataset was divided equally between ChatGPT and Gemini, with images from the same location assigned to the same model to enforce locality-consistent partitioning. The training set contains 177,719 images (152,710 pairs) and the test set 14,806 images (10,020 pairs), each augmented with weather data and text annotations. Of these, 93,834 training images and 14,322 validation images also contain emissions data, enabling end-to-end analysis of atmospheric factors.
@@
Before conducting the model evaluation, the test dataset underwent a thorough manual inspection to ensure the reliability and correctness of the generated annotations of both GPT-4 and Gemini 2.0 models. First, annotators (who were also the authors) verified whether the dominant object in each image aligned with the ground truth class label provided in the original dataset, offering an initial layer of validation to the annotation quality. Second, the annotators were asked targeted questions for each sample, aimed at assessing the accuracy of the generated descriptions with respect to key visual elements and prominent objects present in the image. Each annotator was asked to rate each description on a scale of 1 to 5, over three questions. For each description, the points are summed up, and a threshold of 9 points was used to filter annotations that make it to the test set versus the ones that do not, as elaborated in Figure 3. To account for the inherent subjectivity involved in evaluating such open-ended descriptions, the dataset was split among the annotators in a non-overlapping manner, thereby avoiding inter-rater discrepancies.
@@ IV Methodology
Our approach utilizes two techniques, one based on conversational VLMs and the other based on video-based VLMs, each optimized for different functionality necessary for our task. Conversational VLMs are fine-tuned for interactive conversations, remembering image features, and hence enabling dynamic comparison and support for "what-if" situations raised by users, thus making them interactive. On the other hand, Video-based VLMs are optimized to track temporal variation between a pair of images, which facilitates detailed descriptions for each individual image and its variation. We utilized Video-based VLMs for temporal understanding comparison across them and our conversational VLM ChatENV.
Qwen VLM: We train ChatENV on the Qwen2.5-VL-7B-Instruct model, which combines a Vision Transformer (ViT) image encoder together with a large-language model decoder [35]. This model is a multi-modal version of the Qwen 2.5 model, which is an open-source, light-weight model that is easy to finetune, with competitive performance on VQA and reasoning benchmarks. As shown in Figure 1, the ViT, text encoder, and projection layer are left frozen. Fine-tuning is restricted to light-weight Low-Rank Adaptation (LoRA) adapters placed within every attention and feedforward block within the decoder. A linear probe is optionally added for scalar prediction tasks. A training sample is an RGB aerial tile and a user prompt augmented with structured sensor data. Because environmental sensor readings (e.g., temperature, CO) cannot be directly inferred from an image, we provide them explicitly as user input.
@@
Video-Based VLMs: We employ two recent SOTA models: LLaVA-NeXT-Video [36] and Video-LLaVA [37]. LLaVA-NeXT-Video, a video-adapted variant of LLaVA-NeXT, leverages advanced visual reasoning and improved general knowledge to provide accurate frame-level descriptions and reliable temporal comparisons, particularly excelling in zero-shot video understanding scenarios. Similarly, Video-LLaVA effectively integrates temporal visual information with language understanding, achieving superior performance on video-language benchmarks and making it ideal for articulating subtle visual changes.
@@ V Experiment Setup
LoRA We fine-tuned our model using LoRA [38]. LoRA posits that the pre-trained and fine-tuned weights differ by a low-rank difference. Given a pre-trained matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA decomposes the update as $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and $r\ll\min(d,k)$. Here, $W_{0}$ remains fixed, while $A$ and $B$ are trainable. The forward pass becomes $h=W_{0}x+BAx$, scaled by $\alpha/r$, with $\alpha$ controlling the update magnitude. Typically, $A$ is randomly initialized and $B$ is zero-initialized, ensuring $\Delta W$ starts at zero.
@@
(I) ROUGE [39] computes $n$-gram recall between the generated and reference text. We report ROUGE-L, which rewards the longest common subsequence (LCS) and hence fluency and phrase-level overlap. ROUGE-1 and 2 evaluate unigram and bigram recall, reflecting lexical overlap. ROUGE is particularly well-suited for content coverage estimation.
(II) SBERT Similarity uses Sentence-BERT to encode generated and reference text into high-dimensional vectors, compared against each other using cosine similarity [40]. It detects semantic overlap even on rephrased sentences.
(III) BERTScore [41] is computed by comparing generated and reference tokens using contextual embeddings. We use DeBERTa V3 to calculate precision, recall, and F1, which reflect fluency and semantic similarity across different phrasings.
(IV) COMET [42] is a neural metric that has been trained over human judgments. It projects source, hypothesis, and reference onto a common space and makes predictions for quality ratings for adequacy and fluency. COMET is especially robust for generation tasks and correlates strongly with human evaluations.
@@
All models were trained on a 48GB A6000 GPU for one epoch on the entire dataset. Further tests on ChatGPT and Gemini-specific annotations were done separately. For VLMs based on videos, a sample split equal to roughly 10% of the dataset was built separately for ChatGPT and Gemini and used to test the data size effect on the models’ performance.

### v2

@@ Abstract
Understanding environmental changes from remote sensing imagery is vital for climate resilience, urban planning, and ecosystem monitoring. Yet, current vision language models (VLMs) overlook causal signals from environmental sensors, rely on single-source captions prone to stylistic bias, and lack interactive scenario-based reasoning. We present ChatENV, the first interactive VLM that jointly reasons over satellite image pairs and real-world sensor data. Our framework: (i) creates a 177k-image dataset forming 152k temporal pairs across 62 land-use classes in 197 countries with rich sensor metadata (e.g., temperature, PM10, CO); (ii) annotates data using GPT-4o and Gemini 2.0 for stylistic and semantic diversity; and (iii) fine-tunes Qwen-2.5-VL using efficient Low-Rank Adaptation (LoRA) adapters for chat purposes. ChatENV achieves strong performance in temporal and “what-if” reasoning (e.g., BERT-F1 0.902) and rivals or outperforms state-of-the-art temporal models, while supporting interactive scenario-based analysis. This positions ChatENV as a powerful tool for grounded, sensor-aware environmental monitoring. [^1] The annotated data and code are available on github through: https://github.com/HosamGen/ChatENV/
@@ I Introduction
Monitoring and understanding environmental dynamics are top priorities for addressing pressing issues like climate change, urban growth, and habitat degradation. Remote sensing, particularly through the use of aerial and satellite imaging, offers extensible monitoring of Earth’s surface, enabling real-time detection of changes in land cover, vegetation, urban areas, and other environmental indicators [wulder2012landsat, corradino2019lava]. Recently, Vision-Language Models (VLMs) have come into the spotlight for fusing visual and text data, enabling enhanced automation of complex tasks in remote sensing, ranging from change detection, captioning, and multimodal reasoning [liu2024changeagent, zhang2025georsmllm]. In spite of this, several key challenges remain:
(1) Lack of environmental context, as most VLMs rely solely on images and ignore essential sensor data crucial for understanding and observing geographical changes [liu2024changeagent]; (2) Single-source annotation bias, where using a single language model for data annotation limits linguistic diversity and introduces bias [wang2023skyscript]; (3) Limited spatial and semantic diversity, since widely used datasets like LEVIR-CD [chen2020levircd] and RSICap [hu2025rsgpt] lack the geographic scope and object-class variety necessary for robust modeling [lu2017exploring]; and (4) Lack of interactive reasoning, with current VLMs offering no "what-if" interfaces to support planning and hypothetical decision-making [liu2024changeagent].
@@
  • SOTA Multimodal Dataset. We release the largest sensor‑aware temporal satellite captioning dataset we are aware of, containing over 177k satellite images, divided into 152k time-separated image pairs, across 62 object classes and 197 countries. Each pair is accompanied by rich change captions generated by a dual-model pipeline (GPT-4o and Gemini 2.0), adding linguistic variety, context richness, and improved generalization. All test captions are manually reviewed to ensure accuracy, providing a strong benchmark for environmental change detection tasks.
@@ II Related Work
Benchmark Datasets: Several datasets support geospatial vision–language tasks. For captioning, RSICD [lu2017exploring], NWPU-Captions [Cheng2022], and RSICap [hu2025rsgpt] offer image-text pairs at varying scales. RSVQA [Lobry2020] and BigEarthNet-based extensions [Yuan2022] focus on spatial Visual Question Answering (VQA), while LEVIR-CD [Chen2020] and LEVIR-CC [Liu2022] provide Change Detection (CD) labels and natural language descriptions. ChangeCLIP [Dong2024] benchmarks semantic CD, and SpaceNet [spacenet] provides long-term urban imagery but lacks text annotations. However, most existing datasets lack temporal variety, environmental sensor data, or support for open-ended user queries. There remains a need for benchmarks that support contextual understanding, multimodal learning, and interactive reasoning.
Vision-Language Models: Recent VLMs such as Pixtral [pixtral2024], DeepSeek-VL [lu2024deepseekvl], Janus Pro [chen2025janus], Phi-4-Multimodal [abouelenin2025phi4], and LLaMA-3-Vision [grattafiori2024llama3] demonstrate strong performance on tasks like captioning, visual reasoning, and instruction following. These models integrate vision encoders directly into Large Language Models (LLMs) and are trained on large-scale image–text datasets. While not designed for remote sensing, their scalability and reasoning capabilities make them promising candidates for domain adaptation.
VLMs for Remote Sensing: Domain-specific adaptations of VLMs have emerged to tackle geospatial tasks. For example, RSGPT [hu2025rsgpt] fine-tunes Vicuna on RSICap for aerial captions. RS-LLaVA [bazi2024rsllava] and EarthGPT [Bazi2024, Zhang2024] use CLIP/ViT backbones with LLaMA-based LLMs for captioning and VQA, incorporating multi-sensor data. RemoteCLIP and GeoChat [remoteclip, kuckreja2023geochat] build on large image–text corpora but lack temporal modeling. GeoLLaVA [elgendy2024geollava] uses video-based fine-tuning for change detection, though it still lacks annotation diversity and interactivity.
Captioning and VQA in Aerial Imagery: Initial work on remote sensing VQA began with RSVQA [Lobry2020], which used OpenStreetMap for spatial questions, and was extended by [Yuan2022] using BigEarthNet. These efforts rely on synthetic prompts and lack environmental sensor data. In captioning, RSICD introduced a CNN–RNN pipeline, with NWPU-Captions adding attention mechanisms, and more recent work [Bourcier2024] exploring metadata-aware pretraining; however, it is still limited to static imagery.
@@
RSGPT [hu2025rsgpt]
RS-LLaVA [Bazi2024]
EarthGPT [Zhang2024]
RemoteCLIP [remoteclip]
GeoChat [kuckreja2023geochat]
GeoLLaVA [elgendy2024geollava]
Tree-GPT [Du2023]
ChangeCLIP [Dong2024]
CD-VQA [Yuan2022]
TEOChat [irvin2024teochat]
@@
Interactive Reasoning for Environmental Monitoring: Conversational VLMs like, RS-LLaVA, and EarthGPT enable multi-turn Q&A but lack predictive reasoning. Tree-GPT [Du2023] adds geospatial tools for command-based interaction but cannot simulate hypothetical, counterfactual scenarios. Finally, TEOChat [irvin2024teochat] supports scene monitoring by asking questions about objects that change in an image or disasters affecting a location, but it does not incorporate environmental reasoning based on temperature or other meteorological metrics.
@@ III Data Collection and Pre-processing
Construct Temporal Pairs from fMoW Dataset. fMoW RGB dataset [fmow] consists of 363,571363{,}571 train images and 53,04153{,}041 validation images covering years $2002$ to $2017$ globally. With spatial resolution around $0.3$ mm, it provides increased accuracy for fine-grained change detection between $62$ classes [satmae]. We utilized the dataset’s metadata to pair images based on location while maintaining a minimum temporal gap of $12$ months, as preliminary inspection of the dataset indicated that visible and detectable changes in the scenes typically emerge after approximately one year. Upon filtering and removing duplicates, the final training set has $177k$ images, corresponding to $152k$ pairs, and the test set consists of $14k$ images, which correspond to $10k$ pairs. Table II and Figure 4 provide an overview and a visual distribution of the 62 classes over 197 countries, respectively.
@@
We used fMoW Dataset metadata, which contains geospatial locations, country codes, time zones, and timestamps, to fetch weather and atmospheric emissions data through third-party APIs. For weather features, the VisualCrossing API [visualcrossing] fetched a large set of measures, like temperature, dew point, humidity, and more, by using the captured location and timestamp of each image. Likewise, emissions data were collected using Open-Meteo [openmeteo] for European samples and OpenAQ [openaq] for non-European samples, yielding parameters like $pm2.5$, COCO, $NO_{2}$, and Ozone. We focused on features that appeared most frequently and for which our analyses would be most relevant, and cross-referenced country codes from the fMoW dataset against the country name returned by the APIs to ensure temporal alignment and season-specific context. Features with limited relevance or significant missing data were discarded, while those lacking emissions data were retained to preserve category and geographic diversity. This fusion-mapping of environmental metadata and the dataset’s global, multi-year coverage allowed us to capture temporal, climatic, and atmospheric changes with minimal missing information, supporting downstream modeling work.
@@
We produced descriptive annotations to capture the most important features of each image and summarize temporal variations between matched samples. Two SOTA LLMs, GPT-4o and Gemini 2.0 Flash Thinking, were used to extend generalization and increase robustness. All prompts included the semantic category, season and environmental sensor data, following a consistent prompting strategy for each model. The dataset was divided equally between the LLMs, with images from the same location assigned to the same model to enforce locality-consistent partitioning. The training set contains 177,719 images (152,710 pairs) and the test set 14,806 images (10,020 pairs), each augmented with weather data and text annotations. Of these, 93,834 training images and 14,322 validation images also contain emissions data, enabling end-to-end analysis of atmospheric factors.
@@
Before conducting the model evaluation, the test dataset underwent a thorough manual inspection to ensure the reliability and correctness of the generated annotations of both GPT-4o and Gemini 2.0 models. First, annotators (who were also the authors) verified whether the dominant object in each image aligned with the ground truth class label provided in the original dataset, offering an initial layer of validation to the annotation quality. Second, the annotators were asked targeted questions for each sample, aimed at assessing the accuracy of the generated descriptions with respect to key visual elements and prominent objects present in the image. Each annotator was asked to rate each description on a scale of 1 to 5, over three questions. For each description, the points are summed up, and a threshold of 9 points was used to filter annotations that made it to the test set versus those that did not, as elaborated in Figure 3. To account for the inherent subjectivity involved in evaluating such open-ended descriptions, the dataset was split among the annotators in a non-overlapping manner, thereby avoiding inter-rater discrepancies.
@@ IV Methodology
Our approach utilizes two techniques, one based on conversational VLMs and the other based on video-based VLMs, each optimized for different functionality necessary for our task. Conversational VLMs are fine-tuned for interactive conversations, remembering image features, and hence enabling dynamic comparison and support for "what-if" situations raised by users, thus making them interactive. On the other hand, Video-based VLMs are optimized to track temporal changes between image pairs, which facilitates detailed descriptions for each individual image and its variation. We utilized Video-based VLMs for temporal understanding comparison across them and our conversational VLM "ChatENV".
Qwen VLM: We train ChatENV on the Qwen2.5-VL-7B-Instruct model, which combines a Vision Transformer (ViT) image encoder together with a large-language model decoder [Qwen2.5]. This model is a multi-modal version of the Qwen 2.5 model, which is an open-source, light-weight model that is easy to finetune, with competitive performance on VQA and reasoning benchmarks. As shown in Figure 1, the ViT, text encoder, and projection layer are left frozen. Fine-tuning is restricted to light-weight Low-Rank Adaptation (LoRA) adapters placed within every attention and feedforward block within the decoder. A linear probe is optionally added for scalar prediction tasks. A training sample is an RGB aerial tile and a user prompt augmented with structured sensor data. Because environmental sensor readings (e.g., temperature, CO) cannot be directly inferred from an image, we provide them explicitly as user input.
@@
Video-Based VLMs: We employ two recent SOTA models: LLaVA-NeXT-Video [liu2024llavanext] and Video-LLaVA [liu2024videollava]. LLaVA-NeXT-Video, a video-adapted variant of LLaVA-NeXT, leverages advanced visual reasoning and improved general knowledge to provide accurate frame-level descriptions and reliable temporal comparisons, particularly excelling in zero-shot video understanding scenarios. Similarly, Video-LLaVA effectively integrates temporal visual information with language understanding, achieving superior performance on video-language benchmarks and making it ideal for articulating subtle visual changes.
@@ V Experiment Setup
LoRA We fine-tuned our model using LoRA [lora]. LoRA posits that the pre-trained and fine-tuned weights differ by a low-rank difference. Given a pre-trained matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA decomposes the update as $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and $r\ll\min(d,k)$. Here, $W_{0}$ remains fixed, while $A$ and $B$ are trainable. The forward pass becomes $h=W_{0}x+BAx$, scaled by $\alpha/r$, with $\alpha$ controlling the update magnitude. Typically, $A$ is randomly initialized and $B$ is zero-initialized, ensuring $\Delta W$ starts at zero.
@@
(I) ROUGE [lin2004rouge] computes $n$-gram recall between the generated and reference text. We report ROUGE-L, which rewards the longest common subsequence (LCS) and hence fluency and phrase-level overlap. ROUGE-1 and 2 evaluate unigram and bigram recall, reflecting lexical overlap. ROUGE is particularly well-suited for content coverage estimation.
(II) SBERT Similarity uses Sentence-BERT to encode generated and reference text into high-dimensional vectors, compared against each other using cosine similarity [reimers2019sentence]. It detects semantic overlap even on rephrased sentences.
(III) BERTScore [zhang2019bertscore] is computed by comparing generated and reference tokens using contextual embeddings. We use DeBERTa V3 to calculate precision, recall, and F1, which reflect fluency and semantic similarity across different phrases.
(IV) COMET [rei2020comet] is a neural metric that has been trained on human judgment. It projects the source, hypothesis, and reference onto a common space and makes predictions for quality ratings regarding adequacy and fluency. COMET is especially robust for generation tasks and correlates strongly with human evaluations.
@@
All models were trained on a 48GB A6000 GPU for one epoch on the entire dataset. Further tests on ChatGPT and Gemini-specific annotations were done separately. For video-based VLMs, a subset of approximately 10% of the dataset was built separately for ChatGPT and Gemini and used to evaluate how dataset size influences the models’ performance.